# Secure Wireless Communication Using Distributed Coherent Transmission and Spatial Signal Decomposition

### v1

@@ II-A Synchronization and Localization Based on Two-Way Time Transfer
Once the timing offset $\Delta t$ is estimated, node $m$ can adjust its local clock edge or adjust the transmission time of the waveform to compensate for the timing difference. The transmit timestamps are known to the precision of the clock times, which for the SDR s used in this work (Ettus USRP X310) is accurately known to below a picosecond. The reception times $t_{\mathrm{RX}i}$ thus must be estimated with errors on the order of picoseconds to ensure accurate synchronization. This is accomplished using a spectrally-sparse, two-tone signal that yields a near-optimal estimate of the reception time [9057428]. This approach has been demonstrated to achieve a timing precision of $2.26\text{\,}\mathrm{ps}$ with a $40\text{\,}\mathrm{MHz}$ two-tone waveform.
Once the four timestamps are estimated, the total time of flight $T$ between the two systems can also be estimated, from which the relative distance between the two nodes can be determined after calibrating for the static signal processing latency and RF front end transmission line delays $\delta T$ in node $n$. The calibrated time of flight is given by
$T=\frac{1}{2}\left[(t_{\mathrm{RX}m}-t_{\mathrm{TX}n})+(t_{\mathrm{RX}n}-t_{\mathrm{TX}m})\right]-\delta T$ (3)
@@ II-B Frequency-Transfer
Communications signals are susceptible to phase and frequency offset since they can alter the demodulated symbols, While monolithic systems can rely on receive-side techniques like Costas loops, this cannot correct for frequency differences between two simultaneous transmitters, which must therefore be corrected wirelessly between the nodes. In this work we use a wireless frequency transfer technique demonstrated in [8889331, 7218555], and shown in Fig. 2. The primary node generates a two-tone waveform, the frequency separation of which is the desired reference frequency, in this work $10\text{\,}\mathrm{MHz}$. The secondary node receives the two-tone signals which is then input to a self-mixing circuit that demodulates the reference frequency signal. The self-mixing circuit includes a mixer and an attenuator on the LO path of the mixer to ensure equal amplitude of the two input signals. A filter is used to remove any higher-frequency signals, and then amplified. The demodulated frequency references is then input to a phase-locked loop on the secondary node, enabling a wireless frequency lock.
@@ V-A Hardware Design
The block diagram of the hardware is shown in Fig. 5. Each node used Ettus X310 SDR s, which have an analog bandwidth of $160\text{\,}\mathrm{MHz}$ and a sample rate of $200\text{\,}\mathrm{M}\mathrm{S}\mathrm{a}\mathrm{/}\mathrm{s}$. Node 1 used two SDR s, one to transmit and receive the time transfer waveform to Node 2, and the other for beamforming. The second SDR sent a trigger waveform to the receiving node (an oscilloscope) to collect the calibration and beamforming waveforms. The beamforming and time transfer signals were transferred through log-periodic antennas (L-COM HG2458), the trigger signal was transferred through a different log-periodic antenna (L-COM HG7210LP), and the frequency transfer antenna was a standard gain horn antenna (Narda Model 643). Node 2 had a single SDR which performed the same functions as the first SDR on Node 1 and used the same style of antennas for its beamforming and time-transfer systems. The wireless frequency locking used a signal generator (Keysight E8267D) to generate the frequency reference to send from Node 1 to Node 2 using the standard gain horn antenna. Each node had a computer which controlled the SDR s and which communicated via Wi-Fi for the beamforming scheduling between the nodes. An $80\text{\,}\mathrm{G}\mathrm{S}\mathrm{a}\mathrm{/}\mathrm{s}$ (MSO-X 92004A) oscilloscope was used as the receiver for the experiment and used a log-periodic antenna (L-COM HG2458) for receiving the beamformed signals and a log-periodic antenna (L-COM HG7210LP) along with a $2.1\text{\,}\mathrm{GHz}$ cavity filter to capture the trigger signal. The received beamforming signals were processed offline.
@@ V-C Experimental Setup
At each location, 25 measurements were conducted and then the receiver moved approximately $1\text{\,}\mathrm{m}$ to the right towards node 2. The received data was demodulated with a hard decision demodulator. Because of coordination, multipath, and radio signal interference, we limited the search for Carrier Frequency Offset (CFO) compensation to +/- $50\text{\,}\mathrm{kHz}$.
@@ V-D Experimental Results
Experiments were conducted for both traditional beamforming, where the same data is sent from each node, along with the secure spatial decomposition approach. The measurement results are shown in Table VII. The constellation diagrams for a single waveform at each position are shown in Fig. 7 for the secure approach and Fig. 8 for the traditional approach. At broadside both approaches are able to recover the intended data, but the secure approach does have a non-zero, but very low, SER. This can could be caused by environmental factors along with coordination degradation; because each transmitting node is only transmitting half the information there may be certain environmental and coordination effects that impact the data which is overcome in the traditional approach by transmitting the same data from two locations. However, at off-broadside angles the secure approach shows a significant increase in SER, indicating that data recover at these locations would be more challenging. This is demonstrated with symbol errors above 20 percent for positions 2 through 6; whereas, for the traditional beamforming approach, the data can be fully decoded at all locations, not just broadside. This demonstrates that our approach increases the security over a traditional distributed beamforming array. Table VIII shows the average SNR for measurements at the six positions. The SNR is high at all locations, which demonstrates that all symbol errors are due to the secure transmission approach and not due to low SNR. Note also that the coherent gain of the traditional distributed beamforming approach generally yields higher SNR than the secure approach, as expected since the secure approach does not transmit the same data from the two antennas.

### v2

# Secure Wireless Communication Using Coherent Distributed Transmission and Spatial Signal Decomposition

@@ II-A Synchronization and Localization Based on Two-Way Time Transfer
Once the timing offset $\Delta t$ is estimated, node $m$ can adjust its local clock edge or adjust the transmission time of the waveform to compensate for the timing difference. The transmit timestamps are known to the precision of the clock times, which for the SDR s used in this work (Ettus USRP X310) is accurately known to below a picosecond. The reception times $t_{\mathrm{RX}i}$ thus must be estimated with errors on the order of picoseconds to ensure accurate synchronization. This is accomplished using a spectrally-sparse, two-tone signal that yields a near-optimal estimate of the reception time [9057428]. This approach has been demonstrated to achieve a timing precision of $2.26\text{\,}\mathrm{ps}$ with a $40\text{\,}\mathrm{MHz}$ two-tone waveform [9994246].
Once the four timestamps are estimated, the total time of flight $T_{m,n}$ between the two systems can also be estimated, from which the relative distance between the two nodes can be determined after calibrating for the static signal processing latency and RF front end transmission line delays $\delta T_{m,n}$ in node $n$. The calibrated time of flight is given by
$T_{m,n}=\frac{1}{2}\left[(t_{\mathrm{RX}m}-t_{\mathrm{TX}n})+(t_{\mathrm{RX}n}-t_{\mathrm{TX}m})\right]-\delta T_{m,n}$ (3)
@@ II-B Frequency-Transfer
Communications signals are susceptible to phase and frequency offset since they can alter the demodulated symbols, While monolithic systems can rely on receive-side techniques like Costas loops, this cannot correct for frequency differences between two simultaneous transmitters, which must therefore be corrected wirelessly between the transmit nodes. In this work we use a wireless frequency transfer technique demonstrated in [9210120, 7218555], and shown in Fig. 2. The primary node generates a two-tone waveform, the frequency separation of which is the desired reference frequency, in this work $10\text{\,}\mathrm{MHz}$. The secondary node receives the two-tone signals which is then input to a self-mixing circuit that demodulates the reference frequency signal. The self-mixing circuit includes a mixer and an attenuator on the LO path of the mixer to ensure equal amplitude of the two input signals. A filter is used to remove any higher-frequency signals, and then amplified. The demodulated frequency references is then input to a phase-locked loop on the secondary node, enabling a wireless frequency lock.
@@ V-A Hardware Design
The block diagram of the hardware is shown in Fig. 5. Each node used Ettus X310 SDR s, which have an analog bandwidth of $160\text{\,}\mathrm{MHz}$ and a sample rate of $200\text{\,}\mathrm{M}\mathrm{S}\mathrm{a}\mathrm{/}\mathrm{s}$. Node 1 used two SDR s, one to transmit and receive the time transfer waveform to Node 2, and the other for beamforming. The second SDR sent a trigger waveform to the receiving node (an oscilloscope) to collect the calibration and beamforming waveforms. The beamforming and time transfer signals were transferred through log-periodic antennas (L-COM HG2458), the trigger signal was transferred through a different log-periodic antenna (L-COM HG7210LP), and the frequency transfer antenna was a standard gain horn antenna (Narda Model 643). Node 2 had a single SDR which performed the same functions as the first SDR on Node 1 and used the same style of antennas for its beamforming and time-transfer systems. The wireless frequency locking used a signal generator (Keysight E8267D) to generate the frequency reference to send from Node 1 to Node 2 using the standard gain horn antenna. Each node had a computer which controlled the SDR s and which communicated via Wi-Fi for the beamforming scheduling between the nodes. An $80\text{\,}\mathrm{G}\mathrm{S}\mathrm{a}\mathrm{/}\mathrm{s}$ (MSO-X 92004A) oscilloscope sampled at $20\text{\,}\mathrm{G}\mathrm{S}\mathrm{a}\mathrm{/}\mathrm{s}$ was used as the receiver for the experiment and used a log-periodic antenna (L-COM HG2458) for receiving the beamformed signals and a log-periodic antenna (L-COM HG7210LP) along with a $2.1\text{\,}\mathrm{GHz}$ cavity filter to capture the trigger signal. The received beamforming signals were processed offline.
@@ V-C Experimental Setup
At each location, 25 measurements were conducted and then the receiver moved approximately $1\text{\,}\mathrm{m}$ to the right towards node 2. The received data was demodulated with a hard decision demodulator. Because of coordination, multipath, and radio signal interference, we limited the search for Carrier Frequency Offset (CFO) compensation to $\pm$ $50\text{\,}\mathrm{kHz}$.
@@ V-D Experimental Results
Experiments were conducted for both traditional beamforming, where the same data is sent from each node, along with the secure spatial decomposition approach. The measurement results are shown in Table VII. The constellation diagrams for a single waveform at each position are shown in Fig. 7 for the secure approach and Fig. 8 for the traditional approach. At broadside both approaches are able to recover the intended data, but the secure approach does have a non-zero, but very low, SER. This can could be caused by environmental factors along with coordination degradation; because each transmitting node is only transmitting half the information there may be certain environmental and coordination effects that impact the data which is overcome in the traditional approach by transmitting the same data from two locations. However, at off-broadside angles the secure approach shows a significant increase in SER, indicating that data recover at these locations would be more challenging. This is demonstrated with symbol errors above 20% for positions 2 through 6; whereas, for the traditional beamforming approach, the data can be fully decoded at all locations, not just broadside. This demonstrates that our approach increases the security over a traditional distributed beamforming array. Table VIII shows the average SNR for measurements at the six positions. The SNR is high at all locations, which demonstrates that all symbol errors are due to the secure transmission approach and not due to low SNR. Note also that the coherent gain of the traditional distributed beamforming approach generally yields higher SNR than the secure approach, as expected since the secure approach does not transmit the same data from the two antennas.